# Gender Bias in Perception of Human Managers Extends to AI Managers

### v1

@@ Abstract
As AI becomes more embedded in workplaces, it is shifting from a tool for efficiency to an active force in organizational decision-making. Whether due to anthropomorphism or intentional design choices, people often assign human-like qualities – including gender – to AI systems. However, how AI managers are perceived in comparison to human managers and how gender influences these perceptions remains uncertain. To investigate this, we conducted randomized controlled trials (RCTs) where teams of three participants worked together under a randomly assigned manager – either human or AI – who was presented as male, female, or gender-neutral. The manager’s role was to select the best-performing team member for an additional award. Our findings reveal that while participants initially showed no strong preference based on manager type or gender, their perceptions changed significantly after experiencing the award process. As expected, those who received awards rated their managers as more fair, competent, and trustworthy, while those who were not selected viewed them less favorably. However, male managers – both human and AI – were more positively received by awarded participants, whereas female managers, especially female AI managers, faced greater skepticism and negative judgments when they denied rewards. These results suggest that gender bias in leadership extends beyond human managers and towards AI-driven decision-makers. As AI takes on greater managerial roles, understanding and addressing these biases will be crucial for designing fair and effective AI management systems.
Keywords: Gender bias, AI management, AI and Workplace, Anthropomorphism in AI
Introduction
@@
Gender bias often influences how people assess a manager’s fairness, competence, and leadership qualities, and it has been widely studied. Studies consistently reveal that female managers are often perceived less favorably than their male counterparts [13, 14, 15, 16], potentially due to the discrepancy between the traditional female gender role and the leadership role [17, 18]. Women in managerial positions are often held to higher standards of warmth and likability [19], while also being scrutinized more harshly for assertiveness or decisiveness—traits typically valued in male managers [20]. Male managers are frequently perceived as more competent and authoritative by default [21, 22], which can enhance perceptions of their fairness and leadership capability. Conversely, female managers, despite displaying the same level of competence, may face skepticism regarding their authority or be seen as less credible [18]. Past research indicates that this disparity is especially pronounced in workplace evaluations by male peers [23], where female managers are rated significantly lower than their male counterparts, while female peers provide similar evaluations for both male and female managers [23].
@@
The interplay between manager type (human or AI) and perceived gender in influencing fairness perceptions, particularly in competitive scenarios where only a subset of individuals are awarded, remains underexplored. This study aims to examine how the combined effects of manager type, manager gender, self-perceived contribution, and award status influence the perception of managers’ fairness, trustworthiness, competence, and willingness to work with similar managers in future tasks. Fairness is crucial for morale and acceptance; trustworthiness relates to reliability and ethical standing; competence underpins a manager’s authority and efficacy; and willingness to work in future endeavors reflects the sustainability of the collaboration. These dimensions are essential for fostering positive and effective human-AI team interactions.
@@
To explore these dynamics, we conducted online experiments where participants worked in teams under managerial scenarios varying by manager type and gender. This study aims to provide insights into the nuanced ways that both AI and gendered presentations impact human acceptance of AI managers, ultimately contributing to the growing understanding of effective AI integration in workplaces.
Results
Participants based in the United States were recruited for the online experiment. In the pre-treatment survey, participants were asked to express their perception of managers across different types and genders on four dimensions, using a scale of 0 to 10. Here we present results for perceived fairness; results for other dimensions are available in the SI.
The treatment phase consisted of two rounds: an individual problem-solving task and a team-based task. In the team-based round, participants worked in teams of three, with a randomly assigned manager whose identity labels varied in type and gender. The manager was introduced as responsible for selecting the best player to receive an additional award, whereas in reality, one player was selected to receive an extra award at random. In the next step, the participants were asked similar questions regarding their manager, and their responses to the post-treatment survey were compared with their responses to the pre-treatment survey.
In the pre-treatment survey, participants perceived fairness across manager types and genders averaged around 7, with slight variations (Fig. 1A, D). Female participants gave slightly higher ratings to female managers, while male participants viewed male managers as fairer, though the differences between groups were relatively small. The perceived fairness of managerial decisions shows a weak positive correlation between human and AI managers ( = 0.097) but a strong positive correlation between female and male managers ( = 0.653) (See SI). This suggests that fairness perceptions in the pre-treatment phase may be more influenced by manager type, whereas gender differences appear to have a weaker effect in favor of female managers. Similar patterns are observed across other dimensions as detailed in the SI.
Post-treatment comparisons (Fig. 1B, E) show that receiving an award increased perceived fairness across all manager types. Among female participants, the differences were pronounced, with male-AI managers showing the largest increase and female-human managers the smallest. For male participants, the differences were less pronounced, with male-human managers showing the highest increase and female-human managers the lowest. Conversely, not receiving an award (Fig. 1C, F) generally reduced perceived fairness, with the steepest declines observed for female-AI managers across both female and male participants. Female-human managers experienced greater declines in fairness ratings from female participants than male participants, suggesting that unmet pre-existing expectations may negatively shape post-treatment perceptions. The dynamic shifts observed in fairness perceptions before and after the treatment emphasizes the influence of manager type, gender, and award outcomes.
Figure 2 illustrates the factors influencing post-treatment fairness perception in the linear regression model with interaction terms (Eq. Design and Measures) for female and male participants seperately. is shaped by different persistent factors. For female participants, pre-treatment perceptions of the manager’s gender fairness significantly predict (), whereas for male participants, pre-treatment perceptions of the manager’s type play a similar role ().
Demographic variables such as age and education generally show insignificance effects on participants’ perception of their manager’s decision, except for male participants where a postgraduate degree predicts higher (). Relative to the reference category (awarded by a human-male manager), not receiving an award significantly lowers perceived fairness for female-AI managers ( for female participants; for male participants), female-human managers ( for female participants; for male participants), and AI managers of unspecified gender ( for female participants; for male participants).
Interaction effects between participants’ self-perceived contribution, award receipt, and manager type reveal that fairness perceptions are closely tied to an expectation-reward dynamic. Not receiving an award, especially under high self-perceived contribution, significantly reduces fairness perceptions for both human managers ( for female participants; for male participants) and AI managers ( for female participants; for male participants). The reduction in fairness perception for human managers among female participants is greater than that for AI managers and exceeds the reduction observed for either human or AI managers among male participants. This may stem from higher expectations of fairness and empathy in human leadership from female participants. The results we report here are specifically for perceived fairness. Findings for the other dimensions are generally consistent; see Supporting Information for additional details.
Discussion
Being selected for an award increased participants’ positive perceptions of the manager across all manager types. However, the magnitude of these benefits was moderated by manager gender, with male managers experiencing more pronounced gains than female managers. This aligns with pervasive societal stereotypes that link competence and leadership effectiveness with male figures [31, 19, 18]. The reinforcement of positive recognition for male managers, and especially male-AI managers (Fig. 1), could reflect societal tendencies to perceive male-led successes as more legitimate or expected, a bias that some have argued is deeply ingrained in patriarchal organizational structures [42]. Furthermore, algorithmic neutrality and objectivity [43] may intersect here, as male-AI managers might benefit from a dual layer of presumed objectivity: male leadership and AI precision.
On the other hand, negative outcomes notably harmed perceptions of female managers, particularly female-AI managers, highlighting a compounded vulnerability, where skepticism toward AI decision-making intersects with gender biases to yield harsher judgments. Female human managers also faced a notable decline in the assessment of participants, though to a lesser extent, possibly due to the perception that human managers demonstrate greater accountability and empathy [10]. By contrast, male managers exhibited greater resilience to negative outcomes, likely due to societal leniency afforded to male figures, consistent with status expectations theory [19]. This underscores the moderating role of gender in shaping perceptions of managerial decisions, illustrating a deeply embedded double standard where cultural narratives shield male leaders with presumed competence and external locus of control, while female leaders face greater blame for unfavorable outcomes. These findings resonate with broader feminist critiques that women’s roles and societal positions are often disproportionately scrutinized within a male-dominated societal framework [44, 45].
The amplified negative impact on female-AI managers emphasizes the need to address intersectional biases in workplace dynamics, where gender and technology intersect to create compounded disadvantages. Organizations aiming to integrate AI in leadership roles should consider strategies to mitigate such biases, such as transparency in AI decision-making processes and fostering a culture that values fairness and equity. Furthermore, the results point to the importance of aligning performance outcomes with fairness perceptions, as these evaluations can significantly influence trust, collaboration, and organizational morale.
Our study has limitations that highlight avenues for future research. Conducted in a single setting with one specific task, our findings may not generalize to more complex scenarios involving diverse tasks, larger groups, or mixed-agent interactions between humans and AI in the same environment. Future studies could explore the mechanisms driving these biases, such as cognitive biases, stereotypes, and trust dynamics, to better understand how they shape various dimensions of perceptions, including trustworthiness, competence, fairness, and willingness to collaborate. Additionally, broader contextual factors like cultural norms, team dynamics, and prior experiences with AI or human managers, could also be examined to uncover the nuanced drivers of these perceptions. These insights will be essential for designing systems across a range of settings that foster equitable, trustworthy, unbiased, and effective integration of AI in managerial and collaborative tasks.
The emergence of AI in managerial and societal roles presents an opportunity to move beyond outdated frameworks and mindsets, yet it also risks perpetuating or even amplifying existing gender and sexist stereotypes if left unchecked. While significant scrutiny is placed on ensuring AI is human-centric, fair, and explainable, we must also hold ourselves accountable as humans to embody these principles in our behaviors and societal norms. History has shown how prejudice and discrimination linger in our actions and institutions, and these biases could be magnified as AI becomes more embedded in daily life. To prevent this, it is imperative to address and dismantle these stereotypes now—before they shape the interactions between humans and AI in ways that are counterproductive or harmful. Policymakers currently shaping AI regulation must consider these risks and explore bold measures, such as developing AI systems that are gender-neutral, to break free from reinforcing stereotypes. By confronting these issues today, we set the stage for a more equitable future where both humans and AI contribute to a society free from discrimination.
Materials and Methods
@@ Discussion
Participants
Participants were recruited from the Prolific platform (https://www.prolific.com/), selecting U.S.-based adults aged 18 and older for an online experiment exploring perceptions of decision-making in scenarios involving human and AI managers. The final sample included N = 556 participants, with an average age of 37.5 years ( 12.1). The sample consisted of approximately 48.7% female, 48.4% male, and 2.9% identifying as non-binary or other genders. Participants reported various educational backgrounds, ranging from secondary education to postgraduate degrees. (See details in Supporting Information.)
Design and Measures
Figure 3 illustrates a methodological framework of the experiment. This study employed a mixed experimental design to examine how participants’ perceptions of trustworthiness, competence, fairness, and willingness to work with a manager were influenced by primary factors: manager type, manager gender, award outcome, and self-perceived contribution. Participants were randomly assigned to three-person groups, with each group placed in a different manager condition. In these conditions, participants viewed scenarios in which either a human or an AI manager made a selection decision. Within each manager type, the gender of the manager was also manipulated, with the manager identified as either female, male, or gender unspecified to explore potential biases or differential expectations based on perceived gender.
@@
| Contribution | |||
where is the rating an individual gives to themselves and is the rating an individual gives to teammate i. This self-assessed contribution measure offers insight into how individuals’ perceptions of their own contributions, relative to their peers, influence their judgments of the manager.
@@
The model includes pre-treatment fairness perceptions based on manager type (pre-type-fairness) and manager gender (pre-gender-fairness), along with demographic variables such as age, education, and gender. Interaction terms were included to capture the combined effects of awards, manager type, manager gender, and contribution on post-treatment fairness. Specifically, the model accounts for the three-way interaction between awards, manager type, and manager gender, as well as the three-way interaction between awards, manager type, and contribution. The model was applied separately to female and male participants.
@@
Individual problem-solving task. Following the preliminary survey, participants engaged in two rounds of a game designed to foster both individual and team-based problem-solving. The game involved solving a brain teaser by observing pictures and identifying the “robber”. In the first round, they completed the task individually, allowing them to form an initial self-assessment of their performance.
Manager assignment. After completing the task, participants were informed they would join a team with two teammates for the next round. Each team was randomly assigned a manager, either a human or an AI, who would evaluate the team’s performance. The manager’s gender was also manipulated within each manager type, with the manager identified as male, female, or gender unspecified. The manager would ultimately select the “best player” in the team to receive an extra cash award of £0.5.
Team-based task and collaboration. In the second round, participants collaborated in a team-based setting through an online chat interface to complete the same task as in the first round. They could view their own and their teammates’ responses from the previous round on the slider bars, providing context for discussion. Through the chat, participants worked collectively to arrive at a final team response, simulating real-world team dynamics. All teams completed identical tasks, ensuring a consistent basis for comparing perceptions of the managers across groups.
Contribution assessment. After completing the team task, participants rated their own and their teammates’ contributions on a scale from 0 to 10. The chat history remained visible during this rating phase but was disabled for further input, ensuring that participants could not influence each other’s ratings. This contribution assessment allowed for an additional evaluation of how individuals perceived their own and their peers’ contributions within the team.
Manager decision and exit survey. After a waiting time of 20 seconds, participants received the announcements of the “best player” selected by the manager. Finally, participants completed an exit survey that included the study’s primary measures: perceptions of trustworthiness, competence, and fairness regarding the manager’s decision, as well as their willingness to work with a similar manager in the future.

### v2

@@ Abstract
As AI becomes more embedded in workplaces, it is shifting from a tool for efficiency to an active force in organizational decision-making. Whether due to anthropomorphism or intentional design choices, people often assign human-like qualities, including gender, to AI systems. However, how AI managers are perceived in comparison to human managers and how gender influences these perceptions remains uncertain. To investigate this, we conducted randomized controlled trials (RCTs) where teams of three participants worked together under a randomly assigned manager. The manager was either a human or an AI and was presented as male, female, or gender-unspecified. The manager’s role was to select the best-performing team member for an additional award. Our findings reveal that while participants initially showed no strong preference based on manager type or gender, their perceptions changed significantly after experiencing the award process. As expected, those who received awards rated their managers as more trustworthy, competent, fair, and were more willing to work with similar managers in the future, while those who were not selected viewed them less favorably. However, male managers, whether human or AI, were more positively received by awarded participants, whereas female managers, especially female AI managers, faced greater skepticism and negative judgments when they did not give awards. These results suggest that gender bias in leadership extends beyond human managers to include AI-driven decision-makers. As AI assumes more managerial responsibilities, understanding and addressing these biases will be crucial for designing fair and effective AI management systems.
Keywords: Gender bias, AI management, AI and Work, Anthropomorphism
1 Introduction
@@
Gender bias often influences how people assess a manager’s fairness, competence, and leadership qualities, and it has been widely studied. Studies consistently reveal that female managers are often perceived less favorably than their male counterparts [13, 14, 15, 16], potentially due to the discrepancy between the traditional female gender role and the leadership role [17, 18]. Women in managerial positions are often held to higher standards of warmth and likability [19], while also being scrutinized more harshly for assertiveness or decisiveness – traits typically valued in male managers [20]. Male managers are frequently perceived as more competent and authoritative by default [21, 22], which can enhance perceptions of their fairness and leadership capability. Conversely, female managers, despite displaying the same level of competence, may face skepticism regarding their authority or be seen as less credible [18]. Past research indicates that this disparity is especially pronounced in workplace evaluations by male peers [23], where female managers are rated significantly lower than their male counterparts, while female peers provide similar evaluations for both male and female managers [23].
@@
The interplay between manager type (human or AI) and perceived gender, in shaping perceptions of managers, particularly in competitive scenarios where only a subset of individuals are awarded, remains underexplored. While gender bias is a central concern of this research, it is part of a broader investigation into how multiple interacting factors influence managerial perceptions. This study aims to examine how the combined effects of manager type, manager gender, self-perceived contribution, and award status influence the perception of managers’ trustworthiness, competence, fairness, and willingness to work with similar managers in future tasks. Trustworthiness relates to reliability and ethical standing; competence underpins a manager’s authority and efficacy; fairness is crucial for morale and acceptance; and willingness to work in future endeavors reflects the sustainability of the collaboration. These dimensions are essential for fostering positive and effective human-AI team interactions.
@@
To explore these dynamics, we conducted online experiments where participants worked in teams under managerial scenarios varying by manager type and gender. This study provides insights into the nuanced ways that both AI and gendered presentations impact human acceptance of AI managers, ultimately contributing to the growing understanding of effective AI integration in workplaces. Based on prior literature and theoretical frameworks, we propose the following hypotheses:
H1: Participants who receive an award will rate their manager as more trustworthy, competent, fair, and will express greater willingness to work with a similar manager in the future than participants who do not receive an award.
H2: Managers presented as male (both human and AI) will receive more favorable ratings than those presented as female, especially among participants who receive an award.
H3: Managers presented as female (both human and AI) will receive more negative ratings than those presented as male, especially among participants who do not receive an award.
H4: Participants who perceive themselves as high contributors but do not receive an award will rate their manager as less fair, particularly when the manager is human.
H5: There will be significant interaction effects between manager type, manager gender, award outcome, and self-perceived contribution on participants’ ratings of trustworthiness, competence, fairness, and willingness to work with a similar manager in the future.
2 Materials and Methods
@@
2.1 Participants
Participants were recruited from the Prolific platform (https://www.prolific.com/), selecting U.S.-based adults aged 18 and older for an online experiment exploring perceptions of decision-making in scenarios involving human and AI managers. The sample included N = 556 participants, with an average age of 37.4 years (12.2). The sample consisted of approximately 48.7% female, 48.4% male, and 2.9% identifying as non-binary or other genders. Participants reported various educational backgrounds, ranging from secondary education to postgraduate degrees. (See details in Supporting Information Fig. S10.)
2.2 Design and Measures
Figure 1 illustrates the timeline view of the experimental methodology. This study employed a mixed experimental design to examine how participants’ perceptions of trustworthiness, competence, fairness, and willingness to work with a manager were influenced by primary factors: manager type, manager gender, award outcome, and self-perceived contribution. Participants were randomly assigned to three-person groups, with each group placed in a different manager condition. In these conditions, participants viewed scenarios in which either a human or an AI manager made a selection decision. Within each manager type, the gender of the manager was also manipulated, with the manager identified as either female, male, or gender unspecified to explore potential biases or differential expectations based on perceived gender.
@@ 2.2 Design and Measures
| Contribution |
where is the rating an individual gives to themselves and is the rating an individual gives to teammate i. This self-assessed contribution measure captures how individuals perceive their own input relative to their peers. The rationale is that individual perceptions often depend not only on absolute judgments but on how people view themselves in relation to others. By comparing each player’s self-rating with the average rating they assign to their teammates, we quantify the extent to which individuals perceive themselves as contributing more or less than others. This approach draws on theoretical frameworks from social comparison theory [42], which suggests that individuals evaluate themselves by comparing themselves to others, and with research on equity theory [40, 41], where the perceived balance between one’s contributions (inputs) and rewards (outputs) is crucial in shaping social judgments.
@@
The model includes pre-treatment fairness perceptions based on manager type (pre-type-fairness) and manager gender (pre-gender-fairness), along with demographic variables such as age, education, and gender. Interaction terms were included to capture the combined effects of awards, manager type, manager gender, and contribution on post-treatment fairness. Specifically, the model accounts for the three-way interaction between award outcome, manager type, and manager gender, as well as the three-way interaction between awards, manager type, and contribution. The model was applied separately to female and male participants.
@@ 2.3 Procedure
Individual problem-solving task. After completing the preliminary survey, participants began the first round of the game. A countdown timer appeared at the top center of their screen, along with their randomly assigned ID. They were required to individually complete a visual intelligence task (“Find the Robber” ): read the on-screen description, examine the pictures (SI Fig. S1) based on visual cues, and identify the robber from five options. Participants selected the suspect they believed to be the robber using a slider bar within a 2-minute time limit; otherwise, a default value was automatically recorded.
Manager assignment. After completing the task, participants were informed they would join a team with two teammates for the next round (Figs S2–S5). Each team was randomly assigned a manager, either a human or an AI, who would evaluate the team’s performance. The manager’s gender was also manipulated within each manager type, with the manager identified as male, female, or gender unspecified. The manager would ultimately select the “best player” in the team to receive an extra cash award of £0.5.
Team-based task and collaboration. In the second round, participants collaborated in a team-based setting through an online chat interface to complete the same task as in the first round (SI Fig. S6). They could view their own and their teammates’ responses from the previous round on the slider bars, providing context for discussion. Through the chat, participants worked collectively to arrive at a final team response, simulating real-world team dynamics. All teams completed identical tasks, ensuring a consistent basis for comparing perceptions of the managers across groups.
While the Robber task is relatively simple and does not require advanced math or logic, it engages key cognitive and social processes: perceptual discrimination, judgment under uncertainty (no obvious or guaranteed correct answer), time pressure, and social inference (how participants justify their reasoning and assess others’ input). Despite its simplicity compared to real-world workplace tasks, the Robber task provides a controlled and interpretable model for studying reactions to award decisions in ambiguous, team-based settings where credit attribution is inherently uncertain.
Contribution assessment. After completing the team task, participants rated their own and their teammates’ contributions on a scale from 0 to 10 (SI Fig. S7). The chat history remained visible during this rating phase but was disabled for further input, ensuring that participants could not influence each other’s ratings. This contribution assessment allowed for an additional evaluation of how individuals perceived their own and their peers’ contributions within the team.
Manager decision and exit survey. After a waiting time of 20 seconds, participants received the announcements of the “best player” selected by the manager (SI Figs. S8–S9). Finally, participants completed an exit survey that included the study’s primary measures: perceptions of trustworthiness, competence, and fairness regarding the manager’s decision, as well as their willingness to work with a similar manager in the future.
@@
3 Results
3.1 Comparison of pre- and post-treatment perceptions
In the pre-treatment survey, the mean of participants’ perceptions across manager types and genders was around 7, with slight variations (Fig. 2–5 A, D). Participants generally rated human managers more favorably than AI managers. Female participants gave slightly higher ratings to female managers, and male participants gave slightly higher ratings to male managers, though the overall gender differences were relatively small. The perceptions of managerial decisions show a weak positive correlation between human and AI managers but a strong positive correlation between female and male managers (See SI Fig. S11). This suggests that managerial perceptions in the pre-treatment phase may be more influenced by manager type than by manager gender.
Post-treatment comparisons (Fig. 2–5 B, E) show that receiving an award increased perceptions across all manager types. Male managers are consistently rated higher across all dimensions among both female and male participants. Female participants gave the highest ratings for male-AI, especially in fairness perception. Male participants rated male-human managers more favorably, especially in their willingness to work with them again in the future. Female human managers are consistently rated the lowest across all dimensions among both female and male participants.
Conversely, not receiving an award (Fig. 2–5 C, F) generally reduced perceptions of managers, with the steepest declines observed for female-AI managers across both female and male participants, and a statistically significant effect among male participants on all dimensions. The dynamic shifts observed in evaluations before and after the treatment emphasize the influence of manager type, gender, and award outcomes.
3.2 Effects of demographics and pre-treatment perceptions
Figures 6–9 illustrate the factors influencing post-treatment perceptions according to the linear regression model with interaction terms (Eq. 2.2) for female and male participants separately. Post-treatment perceptions are shaped by different persistent factors. For female participants, pre-treatment perceptions of the manager’s gender-related trustworthiness, competence, and fairness significantly predict the corresponding post-treatment perceptions. In contrast, for male participants, pre-treatment perceptions of the manager’s type play a stronger role in shaping the post-treatment perceptions of all four dimensions. Demographic variables such as age and education generally show no significant effect on participants’ perceptions of their manager’s decision, except for male participants, where a postgraduate degree is associated with higher values of post-treatment ratings of managerial competence and fairness.
3.3 Interaction effects: Award Outcome Manager Type Manager Gender
As shown in Figs. 6–9, relative to the reference category (awarded by a male-human manager), not receiving an award significantly lowers perceptions across all dimensions for female-AI managers, female-human managers, and AI managers of unspecified gender, for both female and male participants. Interestingly, for female participants, not receiving an award also significantly lowers perceptions of all dimensions for human managers of unspecified gender. When not awarded by a male-human manager, female participants showed a significant decline in trust, while male participants rated the manager as less competent and expressed lower willingness to work with a similar manager in the future. Detailed statistics are shown in the Supplementary Information Tabs. S1-S8.
3.4 Interaction effects: Award Outcome Manager Type Contribution
Interaction effects between participants’ self-perceived contribution, award outcome, and manager type reveal that perceptions towards managers are closely tied to an expectation-reward dynamic. As shown in Figs. 6–9, not receiving an award, under high self-perceived contribution, significantly reduces perceptions on all four dimensions for both human managers and AI managers, and for both female and male participants.
The consistent reduction in the evaluations of human managers among female participants is greater than that for AI managers. It exceeds the reduction observed for either human or AI managers among male participants. This may stem from higher expectations in human leadership from female participants. See Supplementary Information Tabs. S1-S8 for detailed statistics.
4 Discussion
Being selected for an award increased participants’ positive perceptions of the manager across all manager types. However, the magnitude of these benefits was moderated by manager gender, with male managers experiencing more pronounced gains than female managers. This aligns with pervasive societal stereotypes that link competence and leadership effectiveness with male figures [31, 19, 18]. The reinforcement of positive recognition for male managers, could reflect societal tendencies to perceive male-led successes as more legitimate or expected, a bias that some have argued is deeply ingrained in patriarchal organizational structures [44].
On the other hand, negative outcomes notably harmed perceptions of female managers, particularly female-AI managers, highlighting a compounded vulnerability, where skepticism toward AI decision-making intersects with gender biases to yield harsher judgments. Female human managers also faced a notable decline in the assessment of participants, though to a lesser extent, possibly due to the perception that human managers demonstrate greater accountability and empathy [10]. By contrast, male managers exhibited greater resilience to negative outcomes, likely due to societal leniency afforded to male figures, consistent with status expectations theory [19]. This underscores the moderating role of gender in shaping perceptions of managerial decisions, illustrating a deeply embedded double standard where cultural narratives shield male leaders with presumed competence and an external locus of control, while female leaders face greater blame for unfavorable outcomes. These findings resonate with broader feminist critiques that women’s roles and societal positions are often disproportionately scrutinized within a male-dominated societal framework [45, 46].
The amplified negative impact on female-AI managers emphasizes the need to address intersectional biases in workplace dynamics, where gender and technology intersect to create compounded disadvantages. Organizations aiming to integrate AI in leadership roles should consider strategies to mitigate such biases, such as transparency in AI decision-making processes and fostering a culture that values fairness and equity. Furthermore, the results point to the importance of aligning performance outcomes with how managerial decisions are perceived, as these evaluations can significantly influence trust, collaboration, and organizational morale.
Female participants and male participants showed comparable patterns in managerial perceptions, though differences also emerged. Female participants who perceive themselves as high contributors rated human managers more negatively when not awarded – possibly due to stronger expectations of fairness from human decision-makers. They also rated female managers more harshly than male participants did under non-award conditions, suggesting a more substantial in-group penalty effect, potentially driven by unmet expectations.
Male participants, when not awarded by a male human manager, rated that manager as significantly less competent and reported lower willingness to work with similar managers in the future. This may reflect that shared identity can heighten expectations, making unmet outcomes more disappointing for male participants. Additionally, postgraduate education among male participants predicted more favorable post-treatment ratings of managerial competence and fairness, suggesting that higher education may reduce negative evaluations.
Our study has limitations that highlight avenues for future research. Conducted in a single setting with one specific task, our findings may not generalize to more complex scenarios involving diverse tasks, larger groups, or mixed-agent interactions between humans and AI in the same environment. Future work could expand on our design by exploring tasks across a broader spectrum of complexity, such as logic-based puzzles, creative group challenges, or multi-stage decision processes that require more sustained effort, information sharing, and interdependence among group members. This would help determine whether the observed patterns in perception and bias generalize to more complex and ecologically valid settings.
Additional studies could also explore the mechanisms driving the observed biases, such as cognitive biases, stereotypes, and trust dynamics, to better understand how they shape various dimensions of perceptions, including trustworthiness, competence, fairness, and willingness to collaborate. Additionally, broader contextual factors like cultural norms, team dynamics, and prior experiences with AI or human managers could also be examined to uncover the nuanced drivers of these perceptions. These insights will be essential for designing systems across a range of settings that foster equitable, trustworthy, unbiased, and effective integration of AI in managerial and collaborative tasks.
The emergence of AI in managerial and societal roles presents an opportunity to move beyond outdated frameworks and mindsets, yet it also risks perpetuating or even amplifying existing gender and sexist stereotypes if left unchecked. While significant scrutiny is placed on ensuring AI is human-centric, fair, and explainable, we must also hold ourselves accountable as humans to embody these principles in our behaviors and societal norms. History has shown how prejudice and discrimination linger in our actions and institutions, and these biases could be magnified as AI becomes more embedded in daily life. To prevent this, it is imperative to address and dismantle these stereotypes now – before they shape the interactions between humans and AI in ways that are counterproductive or harmful. Policymakers currently shaping AI regulation must consider these risks and explore bold measures, such as developing AI systems that are gender-neutral, to break free from reinforcing stereotypes. By confronting these issues today, we set the stage for a more equitable future where both humans and AI contribute to a society free from discrimination.
5 Conclusion
Our study shows that award outcomes shape perceptions toward managers, and these effects are significantly moderated by both the manager’s labeled gender and whether the manager is human or AI. Male managers benefited more from positive outcomes, while female managers, particularly female-AI, were disproportionately penalized after unfavorable outcomes. By integrating insights from social psychology, gender studies, and human-AI interaction, our work contributes to the understanding of how longstanding societal biases can manifest in emerging hybrid work environments. It highlights the nuanced ways that gender and technology intersect in shaping perceptions of managerial trust, competence, fairness, and future collaboration.
As AI systems, and particularly “AI Agents”, take on increasingly prominent roles in decision-making and leadership, addressing these compounded biases is essential. Building equitable, trustworthy, and inclusive systems will require not only fair algorithms but also a deeper understanding of how human perceptions, shaped by gender norms and societal bias, interact with algorithmic decision-makers and influence how those systems are interpreted and received.